### v1

# Spatial-Aware Conformal Prediction for Trustworthy Hyperspectral Image Classification

Abstract
Hyperspectral image (HSI) classification involves assigning specific labels to each pixel to identify various land cover categories. Although deep classifiers have shown high predictive accuracy in this field, quantifying their uncertainty remains a significant challenge, which hinders their application in critical contexts. This study first theoretically evaluates the applicability of Conformal Prediction (CP), an emerging technique for uncertainty quantification, in the context of HSI classification. We then propose a conformal procedure that provides HSI classifiers with trustworthy prediction sets, offering coverage guarantees that ensure these sets contain the true labels with a user-specified probability. Building on this foundation, we introduce Spatial-Aware Conformal Prediction (SACP), which incorporates essential spatial information inherent in HSIs by aggregating non-conformity scores of pixels with high spatial correlation. Both theoretical and empirical results demonstrate that SACP outperforms standard CP in HSI classification. The source code is accessible at https://github.com/J4ckLiu/SACP.
Index Terms:
Hyperspectral image classification, uncertainty quantification, conformal prediction
I Introduction
Hyperspectral imaging technology captures spectral information across a broad wavelength range, providing more detailed pixel features than traditional color images [1, 2], which are limited to three channels. Hyperspectral image (HSI) classification is a fundamental task involving the assignment of discrete labels to each pixel to delineate various land cover categories. Despite the high accuracy achieved by deep classifiers in HSI classification [3, 4, 5, 6, 7], the quantification of their uncertainty remains inadequately explored and presents a substantial challenge. Effective uncertainty quantification is imperative due to the diverse and significant practical applications of HSIs, including surveillance, threat detection [8], mineral detection [9], agriculture [10], environmental protection [11], and defense [12], where errors could lead to serious consequences. For instance, in threat detection or mineral exploration, incorrect classifications could result in misidentifying potential threats or valuable resources, which can have severe repercussions. These applications highlight the urgent need to quantify uncertainty in deep HSI classifier decisions and to advance related methodologies.
Various techniques have been developed to estimate and incorporate uncertainty into predictive models, including confidence calibration [13], Monte-Carlo Dropout [14], and Bayesian neural networks [15]. Although these techniques provide valuable insights into decision uncertainty, they lack theoretical guarantees regarding their effectiveness. This highlights the significance of Conformal Prediction (CP) [16, 17, 18]. CP is a statistical framework that can be applied to any predictive model. Instead of predicting a single label based on the output probabilities of a classifier, this framework generates prediction sets that guarantee the inclusion of the true label with a user-specified coverage probability. Thus, it provides a formal performance guarantee, as follows:
$\mathbb{P}(Y_{\text{test}}\in\mathcal{C}(X_{\text{test}})))\geq 1-\alpha,$ (1)
where $Y_{\text{test}}$ is the true label, $\mathcal{C}(X_{\text{test}})$ is the prediction set, and $\alpha$ is the user-specified error rate. The strong theoretical foundation of CP supports its efficacy across a variety of practical applications, ensuring both reliable and trustworthy decision-making. Conformal predictors may generate prediction sets of size 1 for highly confident predictions and larger sets for those with greater uncertainty. Subsequently, human experts can make the final judgments for the predictions sets with larger sizes. Consequently, CP is a highly suitable candidate for quantifying the uncertainty of deep HSI classifiers.
The application of CP to HSI classification remains unexplored, with two primary considerations that need to be addressed for its effective implementation. Firstly, unlike traditional classification tasks, HSI classification models require access to the entire HSI during their training. This overlap could potentially undermine the validity of CP, which typically relies on the exchangeability of calibration and test data. It is therefore imperative to theoretically examine whether the validity of CP is maintained in the context of HSI classification. Secondly, the performance of CP algorithms is dictated by the non-conformity score function once the base classifier is determined. Although many effective non-conformity score functions have been proposed—such as Adaptive Prediction Sets (APS) [19], which generates prediction sets by accumulating descending softmax values, and Regularized Adaptive Prediction Sets (RAPS) [20], which enhances performance through regularization based on class ranks—these score functions generally neglect the critical spatial information inherent in HSIs. This oversight results in suboptimal prediction sets with unnecessarily large sizes. In summary, addressing the following two questions is essential for the effective application of CP in HSI classification:
  • Does the validity of CP persist in the context of HSI classification?
  • How can spatial information be leveraged to enhance the performance of CP in HSI classification?
In this paper, we first undertake a theoretical analysis to confirm the validity of applying CP to HSI classification, even when classifiers are exposed to pixels from both the calibration and test sets during training. Building upon this foundation, we propose a framework for integrating conformal prediction sets into HSI classifiers. We then theoretically demonstrate that incorporating neighborhood information can enhance the statistical efficiency of prediction sets. Motivated by this, we introduce Spatial-Aware Conformal Prediction (SACP), which refines CP by aggregating non-conformity scores of pixels with high spatial correlation. This approach effectively leverages the spatial information inherent in HSIs to improve overall CP performance. Additionally, we provide theoretical demonstration that SACP guarantees coverage.
We conduct extensive experiments on standard HSI classification benchmarks, including Indian Pines, Pavia University, and Salinas. The experimental results indicate that SACP significantly improves the efficiency of various non-conformity score functions in HSI classification, outperforming standard CP while maintaining satisfactory coverage rates across diverse scenarios. Furthermore, SACP demonstrates its effectiveness with both conventional classifiers, such as CNN-based models, and advanced classifiers, including Transformer-based models, highlighting the versatility of our proposed approach.
Our main contributions are summarized as follows:
  • To the best of our knowledge, we are the first to investigate the application of CP to HSI classification. We conduct a theoretical analysis to evaluate the validity of employing CP for this fundamental task.
  • We introduce a framework for equipping HSI classifiers with reliable conformal prediction sets, which ensure the inclusion of true labels with a user-specified probability.
  • We provide a theoretical demonstration that incorporating spatial information can enhance the statistical efficiency of prediction sets. Based on this insight, we propose SACP, a framework that leverages the spatial information inherent in HSIs to improve CP performance. We also theoretically establish that SACP ensures coverage.
  • Extensive experiments validate the effectiveness of our proposed method. The results show that SACP outperforms standard CP in HSI classification by reducing the average size of prediction sets while maintaining satisfactory coverage rates across various scenarios.
II Background
II-A Conformal Prediction
Notation
In this section, we consider the multi-class classification task involving $K$ classes. Let $\mathcal{X}\subset\mathbb{R}^{d}$ represent the input space, and $\mathcal{Y}=\{1,\ldots,K\}$ denote the label space. Let $(\mathit{X},\mathit{Y})$ be a random data pair drawn from the joint distribution $\mathcal{P_{XY}}$. We use $f_{\theta}:\mathcal{X}\rightarrow\mathbb{R}^{K}$ to denote a pre-trained classifier with parameters $\theta$. For the input $\bm{x}$, $f_{\theta,y}(\bm{x})$ represents the $y$-th element of the logits vector $f_{\theta}(\bm{x})$. The conditional probability of class $y$ is approximated by:
$\hat{\pi}_{\theta,y}(\bm{x})=\sigma(f_{\theta}(\bm{x}))_{y}=\frac{e^{f_{\theta,y}(\bm{x})}}{\sum_{i=1}^{\mathit{K}}e^{f_{\theta,i}(\bm{x})}},$ (2)
where $\sigma$ denotes the softmax function and $\hat{\pi}_{\theta}(\bm{x})=(\hat{\pi}_{\theta,1}(\bm{x}),\hat{\pi}_{\theta,2}(\bm{x}),\dots,\hat{\pi}_{\theta,K}(\bm{x}))$ represents the softmax probability distribution. Typically, the classification result is obtained by $\hat{y}=\mathrm{argmax}_{y\in\mathcal{Y}}\hat{\pi}_{\theta,y}(\bm{x})$.
Pipeline of Conformal Prediction
Instead of predicting a single label based on the output probabilities of a classifier, CP constructs a set-valued mapping $\mathcal{C}:\mathcal{X}\rightarrow 2^{\mathcal{Y}}$ that satisfies the following marginal coverage property:
$\mathbb{P}(\mathit{Y}\in\mathcal{C}(\mathit{X}))\geq 1-\alpha,$ (3)
where $\alpha\in(0,1)$ denotes the user-specified error rate and $\mathcal{C}(\mathit{X})$ is the prediction set, which is a subset of $\mathcal{Y}$.
Before deployment, CP starts with a calibration step using a separate calibration set, $\mathcal{D}_{\text{cal}}:=\{(\bm{x}_{i},y_{i})\}_{i=1}^{n}$, drawn independently from the distribution $\mathcal{P}_{\mathcal{XY}}$. Specifically, we compute the non-conformity score $s_{i}=\mathit{S}(\bm{x}_{i},y_{i})$ for each sample in $\mathcal{D}_{\text{cal}}$, which measures how $y_{i}$ “conforms” to the prediction at $\bm{x}_{i}$. The $1-\alpha$ quantile of the scores $\{s_{i}\}_{i=1}^{n}$ is then determined as the threshold $\tau$. Formally, $\tau$ is obtained as follows:
$\tau=\inf\left\{s\mid\frac{\left|\{i:s_{i}\leq s\}\right|}{n}\geq\frac{\lceil(n+1)(1-\alpha)\rceil}{n}\right\}.$ (4)
Once the threshold $\tau$ is determined, the prediction set for the test instance $\bm{x}_{n+1}$ at the error rate $\alpha$ is generated by:
$\mathcal{C}_{1-\alpha}(\bm{x}_{n+1};\tau):=\{y\in\mathcal{Y}\mid\mathit{S}(\bm{x}_{n+1},y)\leq\tau\},$ (5)
where $\mathcal{C}_{1-\alpha}(\bm{x}_{n+1};\tau)$ includes all labels for which their non-conformity score $\mathit{S}(\bm{x}_{n+1},y)$ does not exceed the threshold $\tau$. This equation illustrates a nesting property of the threshold. Specifically, if $\tau_{1}\leq\tau_{2}$, then
$\mathcal{C}_{1-\alpha}(\bm{x}_{n+1};\tau_{1})\subseteq\mathcal{C}_{1-\alpha}(\bm{x}_{n+1};\tau_{2}).$ (6)
As stated in the following theorem, the prediction sets generated by Eqs. (4) and (5) are guaranteed to achieve finite-sample marginal coverage of the true labels.
Theorem 1.
[16] Let $(\bm{x}_{i},y_{i})_{i=1}^{n}$ and $(\bm{x}_{n+1},y_{n+1})$ be exchangeable data samples. Suppose $\tau$ is calculated by Eq. (4) using $(\bm{x}_{i},y_{i})_{i=1}^{n}$, and the prediction set for $(\bm{x}_{n+1},y_{n+1})$ is given by Eq. (5). Then, the coverage guarantee is as follows:
$\mathbb{P}\left(y_{n+1}\in\mathcal{C}_{1-\alpha}\left(\bm{x}_{n+1};\tau\right)\right)\geq 1-\alpha.$
Non-conformity score function
Once the base classifier is determined, the effectiveness of the resulting prediction set depends critically on the chosen non-conformity score function. Two popular options are Adaptive Prediction Sets (APS) and Regularized Adaptive Prediction Sets (RAPS). APS calculates the non-conformity scores by accumulating the softmax probabilities, sorted in descending order:
$S_{\text{APS}}(\bm{x},y)\!=\!\sum_{y_{i}\neq y}\hat{\pi}_{\theta,y_{i}}(\bm{x})\cdot\mathds{1}_{\{\hat{\pi}_{\theta,y_{i}}(\bm{x})>\hat{\pi}_{\theta,y}(\bm{x})\}}+u\cdot\hat{\pi}_{\theta,y}(\bm{x}),$ (7)
where $u$ is an independent random variable following a uniform distribution on $[0,1]$. However, the softmax probabilities typically exhibit a long-tailed distribution, which facilitates the inclusion of tail classes in the prediction sets. To address this issue, RAPS [20] refines prediction sets by incorporating regularization for the classes. Formally, the non-conformity score function of RAPS is defined as follows:
$\begin{array}[]{r}S_{\textrm{RAPS}}(\bm{x},y)=S_{\textrm{APS}}(\bm{x},y)+\lambda\cdot(o(y,\hat{\pi}_{\theta}(\bm{x}))-k)^{+},\end{array}$ (8)
where $o(y,\hat{\pi}_{\theta}(\bm{x}))$ is the label ranking of $y$, $\lambda$ and $k$ are hyperparameters, and $(z)^{+}$ denotes the positive part of $z$. SAPS [21] reduces the emphasis on probabilities by retaining only the maximum softmax probability and discarding the others. Formally, SAPS is defined as follows:
$S_{\text{SAPS}}(\bm{x},y)=\left\{\begin{array}[]{l}u\cdot\hat{\pi}_{\theta,\max}(\bm{x}),\quad\text{if }o(y,\hat{\pi}_{\theta}(\bm{x}))=1,\\ \hat{\pi}_{\theta,\max}(\bm{x})+(o(y,\hat{\pi}_{\theta}(\bm{x}))-2+u)\cdot\lambda,\text{else},\end{array}\right.$ (9)
where $\lambda$ balances the weight of ranking information, $\hat{\pi}_{\theta,\max}(\bm{x})$ denotes the maximum softmax probability of instance $\bm{x}$, and $u$ is a random variable uniformly distributed on $[0,1]$. It can be concluded that APS, RAPS, and SAPS depend exclusively on softmax probabilities and label ranks.
Evaluation Metric
We evaluate the generated prediction sets using three key metrics: marginal coverage rate (Coverage), average size of the prediction sets (Size), and Size-Stratified Coverage Violation (SSCV). Coverage measures the percentage of samples whose sets contain true labels:
$\text{Coverage}=\frac{1}{|\mathcal{D}_{\text{test}}|}\sum_{(\bm{x},y)\in\mathcal{D}_{\text{test}}}\mathds{1}_{\{y\in\mathcal{C}(\bm{x})\}},$ (10)
which should ideally match $1-\alpha$. Size measures the average length of prediction sets. When coverage is satisfied, a smaller size indicates higher statistical efficiency and is more desirable. Formally, Size is defined as:
$\text{Size}=\frac{1}{|\mathcal{D}_{\text{test}}|}\sum_{(\bm{x},y)\in\mathcal{D}_{\text{test}}}|\mathcal{C}(\bm{x})|.$ (11)
SSCV [20] effectively reflects the conditional coverage rate and is formally defined as:
$\text{SSCV}=100\times\sup_{j}\left|(1-\alpha)-\frac{\left|\{i:y_{i}\in\mathcal{C}(\bm{x}_{i})\text{ and }i\in\mathcal{J}_{j}\}\right|}{|\mathcal{J}_{j}|}\right|,$ (12)
where $\mathcal{J}_{j}$ denotes the partitioned sets, with prediction sets categorized according to their sizes. This metric quantifies the maximum deviation between the observed coverage rate and $1-\alpha$ across the various size categories.
II-B Related Works
Despite the high accuracy achieved by deep classifiers in HSI classification, the quantification of their uncertainty remains inadequately explored and presents a substantial challenge. Uncertainty quantification techniques applied to graph-based HSI classifiers have been investigated in [22]. This study adapts advanced uncertainty quantification models, originally designed for node classification in graphs, to the context of HSI classification. Additionally, prior work [23] utilizes Bayesian deep learning to estimate prediction uncertainty while also enhancing the structural characteristics of the models. However, these techniques lack theoretical guarantees, which underscores the importance of CP [16]. Typically, CP is applied in an inductive manner, also known as split conformal prediction, which employs a separate calibration set to improve computational efficiency. Notable variations include cross-validation methods [16] and the jackknife technique [24]. CP has demonstrated its versatility across various fields, including classification [25, 21], regression [26], and specialized areas such as image generative models [27], robotic control [28], graph neural networks [29, 30], and large language models (LLMs) [31, 32, 33]. Despite its broad application, the use of CP in HSIs remains unexplored, which motivates our research.
III Validity of CP on HSI
HSI Classification
Let $\mathcal{I}\in\mathbb{R}^{H\times W\times U}$ be the HSI to classify, where $H$ and $W$ denote the height and width in pixels, and $U$ denotes the number of spectral bands. Let $M=HW$ be the total number of pixels. Typically, HSI classification aims to predict the label $y\in\mathcal{Y}$ for each pixel in the image, given that a random subset of labels known. The pixels with unknown labels appear randomly at any location. When employing a deep classifier, the $(i,j)$-th pixel is generally considered as the center, with information from the small cuboid $B\in\mathbb{R}^{P\times P\times U}$ utilized to extract the joint spectral-spatial feature representation for the central pixel [34] [^1] In prior research, the terms “cuboids”, “cubes”, and “patches” are often used interchangeably to describe local regions of HSIs. For clarity, this work consistently uses the term “cuboids”. . Here, $P$ denotes the patch size of $B$. Correspondingly, the dataset constructed from $\mathcal{I}$ consists of $M$ such cuboids, as follows:
$\mathcal{D}_{\mathcal{I}}:=\{(B_{i},y_{i})\mid 1\leq i\leq M\},$ (13)
where $y_{i}$ represents the label of the central pixel of $B_{i}$. Here, $\{(B_{i},y_{i})\}_{i=1}^{M}$ are exchangeable to $\mathcal{D}_{\mathcal{I}}$. Typically, the HSI dataset is partitioned into $\mathcal{D}_{\text{train}}$, $\mathcal{D}_{\text{val}}$, and $\mathcal{D}_{\text{test}}$ for training, validation, and evaluation of the deep classifier. Essentially, the HSI classifier is engineered to accurately predict the label of the central pixel within each cuboid in $\mathcal{D}_{\text{test}}$.
Exchangeability and validity of CP on HSI classification
As outlined in Theorem 1, CP necessitates that calibration and test samples be exchangeable. However, HSI classifiers typically operate on pixel cuboids as the fundamental unit of training, rather than individual pixels, and thus require access to the entire HSI during training. Therefore, when partitioning $\mathcal{D}_{\mathcal{I}}$ into $\mathcal{D}_{\text{train}}$, $\mathcal{D}_{\text{cal}}$, and $\mathcal{D}_{\text{test}}$, cuboids in $\mathcal{D}_{\text{train}}$ may overlap with those in $\mathcal{D}_{\text{cal}}$ or $\mathcal{D}_{\text{test}}$ due to shared pixels, which potential breaks the validity of CP. In this section, we demonstrate that, under a general permutation-invariance condition, the exchangeability of the non-conformity scores remains valid even when both calibration and test data are used during training. This theoretical analysis conclusively establishes the validity of applying CP to HSI classification.
Assumption 1.
For any instance cuboid $B$ and any label $y\in\mathcal{Y}$, the non-conformity score $S(B,y)$ is invariant under any permutation of the union $\mathcal{D}_{\text{cal}}\cup\mathcal{D}_{\text{test}}$.
Assumption 1 implies that the non-conformity scores for any instance-label pairs should remain unaffected by the choice of $\mathcal{D}_{\text{cal}}$, thereby imposing a permutation-invariance condition on the training of the HSI classifier. This ensures that the output probabilities and, consequently, the non-conformity scores are invariant to permutations of the calibration and test data of the HSI. HSI classifiers typically satisfy this assumption, as they do not take into account the ordering of pixels (or cuboids) during training. Based on this fact, we then demonstrate that any HSI classifier can produce valid prediction sets and offer a test-time coverage guarantee.
Theorem 2.
Let $\{(B_{i},y_{i})\}_{i=1}^{n}$ and $\{(B_{n+j},y_{n+j})\}_{j=1}^{m}$ denote the calibration and test sets, respectively. Assume that the non-conformity score function $S$, derived from the pre-trained HSI classifier, satisfies Assumption 1. Define the non-conformity scores for the calibration and test sets as $\{s_{i}\}_{i=1}^{n}$ and $\{s_{n+j}\}_{j=1}^{m}$, where $s_{i}:=S(B_{i},y_{i})$. Under these conditions, the scores $\{s_{i}\}_{i=1}^{n}$ and $\{s_{n+j}\}_{j=1}^{m}$ are exchangeable.
The proof of Theorem 2 is detailed in Appendix A. This theorem demonstrates that exchangeability is preserved and establishes the validity of applying CP to HSI classification. We also empirically demonstrate the coverage guarantee in our experiments, as detailed in Table II. A similar analysis is also discussed in prior work [30], which establishes the applicability of CP to Graph Neural Networks. Building on this theoretical foundation, we propose a framework to equip HSI classifiers with reliable conformal prediction sets.
Conformal HSI Classification
For an HSI $\mathcal{I}$ with a random subset of labels known, the following steps outline how to generate prediction sets for the remaining unlabeled pixels to include their true labels with a coverage rate of $\alpha$:
  • 1.
    For each labeled pixel, construct a cuboid and randomly partition these cuboids into two subsets: $\mathcal{D}_{\text{train}}$, containing $N_{1}$ samples, and $\mathcal{D}_{\text{cal}}$, containing $N_{2}$ samples. Similarly, construct a cuboid for each unlabeled pixel to form the set $\mathcal{D}_{\text{test}}$, which includes $N_{3}$ instances.
  • 2.
    Train a HSI classifier $\hat{\pi}_{\theta}$ using the training subset $\mathcal{D}_{\text{train}}$.
  • 3.
    Define the score function $S$ based on the classifier $\hat{\pi}_{\theta}$ and compute the non-conformity scores $s_{i}=S(B_{i},y_{i})$ for each sample in the calibration set $\mathcal{D}_{\text{cal}}$.
  • 4.
    Determine the threshold $\tau$ using $\{s_{i}\}_{i=1}^{N_{2}}$ by solving:
    $\tau=\inf\left\{s\mid\frac{\left|\{i\mid s_{i}\leq s\}\right|}{N_{2}}\geq\frac{\lceil(N_{2}+1)(1-\alpha)\rceil}{N_{2}}\right\}.$
  • 5.
    For each $B_{i}\in\mathcal{D}_{\text{test}}$, derive the prediction set as follows:
    $\mathcal{C}_{1-\alpha}(B_{i};\tau):=\{y\in\mathcal{Y}\mid S(B_{i},y)\leq\tau\}.$
Theorem 2 implies that the non-conformity scores of $\mathcal{D}_{\text{cal}}$ and $\mathcal{D}_{\text{test}}$ are exchangeable. Consequently, it can be inferred that the prediction set $\mathcal{C}_{1-\alpha}(B_{i};\tau)$ will include the true label of the central pixel of $B_{i}$ with a probability of $1-\alpha$. For the non-conformity score function $S$, we can employ APS, RAPS or SAPS. However, these score functions depend exclusively on softmax probabilities and label ranks, neglecting the crucial spatial information inherent in HSIs. This observation motivates the exploration of integrating spatial information into the score calculation to enhance CP performance.
Figure 1: Overview of SACP: Raw non-conformity scores lead to excessively large prediction sets. SACP addresses this issue by refining the scores through aggregation of instances with high spatial correlation. This approach is a simple yet effective solution for enhancing the efficiency of prediction sets.
IV Spatial-Aware Conformal Prediction
IV-A Motivation
Numerous studies have demonstrated that leveraging spatial features can significantly enhance HSI classification accuracy [35, 36]. However, in the context of CP, existing non-conformity scores are computed solely based on the output conditional probability of the central pixel, which results in a considerable neglect of spatial information. Below, we present an analysis that illustrates the impact of incorporating spatial information into the calculation of non-conformity scores.
Given a non-conformity score function $S$ based on an HSI classifier, the score for a pair $(B_{i},y_{i})$ is $S(B_{i},y_{i})$. To incorporate information from neighboring cuboids, we define the adjusted score function $\tilde{S}(B_{i},y_{i})$ as follows:
$\tilde{S}(B_{i},y_{i})=S(B_{i},y_{i})+\frac{1}{|\mathcal{N}_{i}|}\sum_{B_{j}\in\mathcal{N}_{i}}S(B_{j},y_{i}),$ (14)
where $\mathcal{N}_{i}$ denotes the set of neighboring cuboids around the central pixel of $B_{i}$. Intuitively speaking, since neighboring pixels in HSI often belong to the same class, if $y_{i}$ is indeed the true label for $B_{i}$, it is likely that $y_{i}$ is also the true label for $B_{j}\in\mathcal{N}_{i}$. Consequently, the term $\frac{1}{|\mathcal{N}_{i}|}\sum_{B_{j}\in\mathcal{N}_{i}}S(B_{j},y_{i})$ will generally be small. Conversely, if $y_{i}$ is not the correct label for $B_{i}$, this summation is expected to be larger. Thus, the adjusted score $\tilde{S}(B_{i},y_{i})$ will increase less for correct instance-label pairs and more for incorrect ones. The advantages of this adjustment are further elucidated by the following proposition.
Proposition 1.
Consider two different score functions $S$ and $\hat{S}$ based on an HSI classifier. Let $\mathcal{D}_{\text{cal}}$ represent the calibration set, containing $n$ samples, and let $\mathcal{D}_{\text{test}}$ denote the test set, containing $n^{\prime}$ instances. We define $\hat{\mathcal{D}}_{\text{test}}$ as follows:
$\hat{\mathcal{D}}_{\text{test}}:=\{(B_{i},y)\mid(B_{i},y_{i})\in\mathcal{D}_{\text{test}},y\in\mathcal{Y}\}.$
Let $\{s_{i}\}_{i=1}^{n}$ and $\{\hat{s}_{i}\}_{i=1}^{n}$ be the respective scores for $\mathcal{D}_{\text{cal}}$ based on $S$ and $\hat{S}$. The scores for $\hat{\mathcal{D}}_{\text{test}}$ are denoted by $\{t_{k}\}_{k=1}^{n^{\prime}|\mathcal{Y}|}$ and $\{\hat{t}_{k}\}_{k=1}^{n^{\prime}|\mathcal{Y}|}$, respectively. We define
${R}(\{s_{i}\}_{i=1}^{n},\{t_{k}\}_{k=1}^{n^{\prime}|\mathcal{Y}|})=\frac{1}{n\times n^{\prime}|\mathcal{Y}|}\sum_{i=1}^{n}\sum_{k=1}^{n^{\prime}|\mathcal{Y}|}\mathds{1}_{\{s_{i}>t_{k}\}},$
which represents the empirical probability that a score from $\mathcal{D}_{\text{cal}}$ is greater than the one from $\hat{\mathcal{D}}_{\text{test}}$. Then, we have
$\phantom{~{}~{}~{}~{}}{R}(\{s_{i}\}_{i=1}^{n},\{t_{k}\}_{k=1}^{n^{\prime}|\mathcal{Y}|})$ $>{R}(\{\hat{{s}}_{i}\}_{i=1}^{n},\{\hat{t}_{k}\}_{k=1}^{n^{\prime}|\mathcal{Y}|})\Leftrightarrow$
$\sum_{B:(B,y)\in\mathcal{D}_{\text{test}}}\int_{0}^{1}|\mathcal{C}_{1-\alpha}(B)|\,\mathrm{d}\alpha$ $>\sum_{B:(B,y)\in\mathcal{D}_{\text{test}}}\int_{0}^{1}|\hat{\mathcal{C}}_{1-\alpha}(B)|\,\mathrm{d}\alpha,$
where $\mathcal{C}(\cdot)$ and $\hat{\mathcal{C}}(\cdot)$ represent the prediction sets yielded by $S$ and $\hat{S}$, and $\alpha$ denotes the error rate.
The proof of Proposition 1 is provided in Appendix B. It suggests that the probability of a randomly selected score from $\mathcal{D}_{\text{cal}}$ exceeding a randomly selected one from $\hat{\mathcal{D}}_{\text{test}}$ serves as an indicator of CP performance: a lower probability indicates more efficient prediction sets. For the adjusted score $\tilde{S}$ in Eq. (14), which incorporates aggregation within neighborhoods, there is typically a larger increase in scores for $\hat{\mathcal{D}}_{\text{test}}$ compared to those in $\mathcal{D}_{\text{cal}}$. This discrepancy arises because the instance-label pairs in $\mathcal{D}_{\text{cal}}$ are always correctly matched, whereas $\hat{\mathcal{D}}_{\text{test}}$ predominantly contains mismatched ones. Thus, compared to standard CP, the incorporation of spatial information (via Eq. (14)) is expected to reduce the probability defined in Proposition 1, resulting in more efficient prediction sets.
IV-B Methodology
In our preceding analysis, we demonstrate that incorporating scores from neighboring pixels potentially enhances the efficiency of prediction sets. To fully leverage the spatial information inherent in HSIs, we introduce Spatial-Aware Conformal Prediction (SACP). Central to this framework is the Score Aggregation Operator $\mathcal{V}_{k}$, which is defined as follows:
$\mathcal{V}_{k}(B_{i},y)=(1-\lambda)\mathcal{V}_{k-1}(B_{i},y)+\frac{\lambda}{|\mathcal{N}_{i}|}\sum_{B_{j}\in\mathcal{N}_{i}}\mathcal{V}_{k-1}(B_{j},y),$ (15)
where $k$ represents the iteration number, and $\mathcal{V}_{0}=S$ is the utilized base score function. The parameter $\lambda$ serves as a trade-off parameter to balance the influence of the neighborhood information. The term $\mathcal{N}_{i}$ denotes the set of neighboring cuboids around the central pixel of $B_{i}$ that are not in $\mathcal{D}_{\text{train}}$. For $k=1$, each score is aggregated with its immediate neighbors. As $k$ increases, the scores are computed based on previously aggregated scores, thus incorporating spatial information from progressively larger neighborhoods. An overview of SACP is illustrated in Figure 1. While raw scores often result in excessively large prediction sets, SACP refines these scores by aggregating them according to spatial correlations, thereby yielding more efficient prediction sets. Additionally, the complete pipeline of SACP is detailed in Algorithm 1.
Algorithm 1 Spatial-Aware Conformal Prediction (SACP)
1:  Input: Training set $\mathcal{D}_{\text{train}}$, calibration set $\mathcal{D}_{\text{cal}}$, test cuboid $B_{n+1}$, error rate $\alpha$, and parameters $k$ and $\lambda$.
2:  Output: Conformal prediction set $\mathcal{C}_{1-\alpha}(B_{n+1};\hat{\tau})$.
3:  Step 1: Train a deep HSI classifier $\hat{\pi}_{\theta}$ using $\mathcal{D}_{\text{train}}$.
4:  Step 2: Define the score function $S$ based on $\hat{\pi}_{\theta}$.
5:  Step 3: Compute the scores for $\mathcal{D}_{\text{cal}}$ and aggregate the results using the Score Aggregation Operator defined in Eq. (15). Denote the aggregated scores as $\{\hat{s}_{i}\}_{i=1}^{{\scriptstyle|\mathcal{D}_{\text{cal}}|}}$.
6:  Step 4: Determine the threshold $\hat{\tau}$ by:
$\hat{\tau}=\inf\left\{s\mid\frac{\left|\{i\mid\hat{s}_{i}\leq s\}\right|}{N_{2}}\geq\frac{\lceil(N_{2}+1)(1-\alpha)\rceil}{N_{2}}\right\}.$
7:  Step 5: For $B_{n+1}$, the prediction set is derived as:
$\hat{\mathcal{C}}_{1-\alpha}(B_{n+1};\hat{\tau}):=\{y\in\mathcal{Y}\mid\mathcal{V}_{k}(B_{n+1},y)\leq\hat{\tau}\},$
where $\mathcal{V}_{k}$ is the Score Aggregation Operator in Eq. (15).
To demonstrate that SACP can produce trustworthy prediction sets, we also theoretically establish that SACP ensures coverage, which is formalized in the following proposition:
Proposition 2.
Consider the Score Aggregation Operator $\mathcal{V}_{k}$, with $\mathcal{V}_{0}=S$ as the initial score function. Let $\{(B_{i},y_{i})\}_{i=1}^{n}$ be the calibration set and $(B_{n+1},y_{n+1})$ be a sample from the test set. Denote the aggregated scores for the calibration set by $\{\hat{s}_{i}\}_{i=1}^{n}$, where $\hat{s}_{i}=\hat{s}_{i}^{(k)}:=\mathcal{V}_{k}(B_{i},y_{i})$. Suppose the threshold $\hat{\tau}$ is computed by:
$\hat{\tau}=\inf\left\{s\mid\frac{\left|\{i\mid\hat{s}_{i}\leq s\}\right|}{n}\geq\frac{\lceil(n+1)(1-\alpha)\rceil}{n}\right\}.$
Then, the following coverage guarantee holds:
$\mathbb{P}\left(y_{n+1}\in\hat{\mathcal{C}}_{1-\alpha}(B_{n+1};\hat{\tau})\right)\geq 1-\alpha,$
where $\hat{\mathcal{C}}_{1-\alpha}(B_{n+1};\hat{\tau})$ is formulated as
$\hat{\mathcal{C}}_{1-\alpha}(B_{n+1};\hat{\tau}):=\{y\in\mathcal{Y}\mid\mathcal{V}_{k}(B_{n+1},y)\leq\hat{\tau}\}.$
Proposition 2 can be proved using mathematical induction, and the proof is provided in Appendix C. This proposition formally establishes that SACP maintains the coverage guarantee and can generate prediction sets with the specified coverage probability. Moreover, it offers several compelling advantages:
  • Model-agnostic: SACP is a post-hoc approach that can be integrated into any pre-trained HSI classifier without requiring modifications to the model itself.
  • Easy to use: SACP requires minimal hyperparameter tuning, as it exhibits insensitivity to the iteration count $k$ and the term weight $\lambda$ (see Figure 3).
  • Computationally efficient: SACP maintains computational efficiency comparable to that of standard CP (see the discussion in Section V-C).
V Experiments
V-A Experimental Setup
Datasets
We utilize the commonly used benchmarks Indian Pines (IP), Pavia University (PU), and Salinas (SA) datasets for our evaluations [^2] These datasets are available at https://www.ehu.eus/ccwintco/index.php/Hyperspectral_Remote_Sensing_Scenes. Specifically, IP consists of $145\times 145$ pixels and 200 valid spectral bands, representing 16 different land-cover classes. PU includes $610\times 340$ pixels with 103 usable spectral bands, covering 9 land-cover categories. SA comprises $512\times 217$ pixels with 204 effective spectral bands, corresponding to 16 ground-truth classes. For all these HSI datasets, we first perform random sampling of HSI cuboids to create the training set. The remaining data is then arbitrarily divided into calibration and test sets.
TABLE I: Allocation of the three datasets across different classifiers.
Model $\mathcal{D}_{\mathcal{I}}$ $|\mathcal{D}_{\text{train}}|$ $|\mathcal{D}_{\text{cal}}|$ $|\mathcal{D}_{\text{test}}|$
1D-CNN IP 250 4996 5003
PU 103 21336 21337
SA 244 26943 26942
3D-CNN IP 128 5057 5064
PU 52 21362 21362
SA 122 27003 27004
HybridSN IP 128 5057 5064
PU 52 21362 21362
SA 122 27003 27004
SSTN IP 128 5057 5064
PU 52 21362 21362
SA 122 27003 27004
Models
To validate the generalizability of SACP, we employ four widely-used HSI classifiers, incorporating both CNN-based and Transformer-based models with various scales and architectures. For CNN-based models, we select 1D-CNN [37], 3D-CNN [38], and HybridSN [39]. The 1D-CNN and 3D-CNN differ in the dimensions of their convolutional kernels, while HybridSN utilizes multi-dimensional kernels to enhance feature extraction. For the Transformer-based model, we use SSTN [40]. All HSI classifiers are initially pre-trained on the three datasets and then evaluated on the corresponding test data. Table I details the sizes of the training, calibration, and test datasets for each dataset across different models. Due to its relatively poor performance, the 1D-CNN is allocated a larger amount of training data compared to the other three models. For clarity, Table II demonstrates the Overall Accuracy (OA) and Average Accuracy (AA) results for the pre-trained HSI classifiers across the three datasets. The definitions of OA and AA are given as follows:
$\text{OA}=\frac{1}{|\mathcal{D}_{\text{test}}|}\sum\limits_{i=1}^{|\mathcal{D}_{\text{test}}|}\mathds{1}_{\{\hat{y}_{i}=y_{i}\}},$ (16)
$\text{AA}=\frac{1}{|\mathcal{Y}|}\sum\limits_{y_{j}\in\mathcal{Y}}\frac{\sum_{i=1}^{|\mathcal{D}_{\text{test}}|}\mathds{1}_{\{\hat{y}_{i}=y_{i},y_{i}=y_{j}\}}}{\sum_{i=1}^{|\mathcal{D}_{\text{test}}|}\mathds{1}_{\{y_{i}=y_{j}\}}},$ (17)
where $\mathcal{D}_{\text{test}}:=\{(B_{i},y_{i})\}_{i=1}^{|\mathcal{D}_{\text{test}}|}$, $\hat{y}_{i}=\mathrm{argmax}(\hat{\pi}_{\theta}(B_{i}))$, $\mathds{1}$ is the indicator function, and $j\in\{1,\dots,|\mathcal{Y}|\}$. Note: In HSI classifier pre-training, the primary focus is not on maximizing their performance metrics such as OA and AA, but rather on achieving model diversity. Our goal is to validate the effectiveness—i.e., the generalizability—of our approach across different model architectures and training processes, irrespective of their performance levels. Given that our primary goal is not to optimize classification performance, we utilize only a minimal amount of data for training these HSI classifiers.
TABLE II: The results of OA and AA on various datasets of different classifiers.
$\mathcal{D}_{\mathcal{I}}$ Metrics 1D-CNN 3D-CNN HybridSN SSTN
IP OA (%) 68.44 68.93 75.25 88.35
AA (%) 64.38 69.10 74.53 87.43
PU OA (%) 70.45 68.97 68.31 89.39
AA (%) 67.88 65.81 51.82 88.26
SA OA (%) 86.44 86.11 87.91 88.52
AA (%) 85.20 87.61 87.35 88.61
TABLE III: Performance comparison of prediction sets generated by SACP and standard CP across various HSI classifiers and datasets with different error rates $\alpha$. Bold numbers indicate that SACP is superior. $\downarrow$ indicates that smaller values are better.
[…]
$\alpha=0.05$ $\alpha=0.1$
Coverage Size ($\downarrow$) SSCV ($\downarrow$) Coverage Size ($\downarrow$) SSCV ($\downarrow$)
[…]
Implementation Details
The training configurations for all four HSI classifiers adhere to their recommended settings. Each model is trained for 200 epochs utilizing the Adam optimizer [41], with a batch size of 128 and a learning rate of 0.002. For SACP, the trade-off parameter $\lambda$ is fixed at 0.5, and the iteration count $k$ is set to 1 for all experiments. The primary evaluation metrics are Coverage, Size, and SSCV, as detailed in Section II-A. To enhance statistical stability, each experiment is repeated 30 times, and the results are averaged for reporting. All experiments are performed on an NVIDIA GeForce RTX 3090 using PyTorch [42].
V-B Main Results
SACP yields efficient prediction sets
We compare the performance of prediction sets generated by standard CP and SACP across various HSI classifiers and datasets with different error rates $\alpha$. The results, presented in Table III, demonstrate that SACP consistently maintains the desired coverage of $1-\alpha$, empirically validating its coverage guarantee as outlined in Proposition 2. In other words, for each pixel, the true labels are reliably included in their prediction set with a user-specified probability. Formally, this is expressed as:
$\mathbb{P}(Y_{\text{test}}\in\mathcal{C}(X_{\text{test}})))\geq 1-\alpha.$ (18)
Additionally, SACP consistently produces more efficient prediction sets while maintaining a satisfactory conditional coverage rate. For instance, when applying 3D-CNN to the IP dataset, SACP reduces the Size from 5.73 for APS to 3.27, representing a 42.9% decrease. Similarly, for RAPS on the same dataset with 1D-CNN, SACP lowers the Size from 3.64 to 2.26, reflecting a 37.9% reduction. These improvements are observed across different error rates, non-conformity score functions, model architectures, and datasets, highlighting the robustness and generalizability of SACP. It is important to note that, unlike traditional classification tasks, SAPS and RAPS do not consistently outperform APS in the context of HSI classification. This further underscores the challenges associated with CP in HSI classification. Nevertheless, SACP consistently enhances the performance of various non-conformity score functions, affirming its effectiveness.
SACP consistently reduces prediction set size across various pixels
To further demonstrate the effectiveness of SACP, we visualize the size of prediction sets for each pixel in $\mathcal{D}_{\text{cal}}$ and $\mathcal{D}_{\text{test}}$ in Figure 2. This experiment is conducted on the IP and PU datasets using SSTN, with APS as the base score function and $\alpha$ set to 0.05. In the visualization, brighter pixels represent larger prediction sets. The results clearly indicate that SACP significantly reduces the size of prediction sets compared to standard CP. Moreover, standard CP often produces unusually large prediction sets, even for interior pixels, as evidenced in the lower-left corner of Figure 2(d). Typically, edge pixels, which are more challenging to predict, tend to have larger prediction sets, while interior pixels have smaller sets. SACP effectively addresses this issue by consistently reducing prediction set sizes for both edge and interior pixels. This improvement underscores the superior performance of our approach in conformal HSI classification.
Figure 2: Visualizations of the average size of prediction sets for each pixel in $\mathcal{D}_{\text{cal}}$ and $\mathcal{D}_{\text{test}}$ are presented for the (a) IP and (e) PU datasets. The true labels for each class are shown in (b) and (f). The brightness of each pixel indicates the size of its prediction set, with brighter pixels representing larger sets. Panels (c) and (g) display the set sizes with SACP, while panels (d) and (h) show the set sizes with standard CP.
SACP is insensitive to hyperparameters
In SACP, two hyperparameters are crucial: $\lambda$ and $k$. The parameter $\lambda$ serves as a trade-off, balancing the influence of the neighborhood, while $k$ specifies the number of iterations. Increasing $k$ allows SACP to incorporate spatial information from progressively larger neighborhoods. This section examines the impact of variations in $\lambda$ and $k$ on the performance of SACP. The experiment is conducted on all selected datasets using SSTN, with APS as the score function and $\alpha$ set to 0.05. The results are illustrated in Figure 3. Figure 3a demonstrates that SACP consistently outperforms the baseline across different values of $\lambda$, with the size of prediction sets decreasing as $\lambda$ increases. Figure 3b illustrates that SACP consistently achieves superior performance as the number of iterations $k$ changes. Overall, the results suggest that SACP is relatively insensitive to variations in $\lambda$ and $k$, consistently surpassing the baseline. For simplicity, we employ a moderate range of hyperparameters, consistently setting $\lambda=0.5$ and $k=1$ throughout the experiments.
(a) 1
(b) 2
Figure 3: Effect of $\lambda$ and $k$ on Prediction Set Size: Each subplot presents results from three comparative experiments conducted on different datasets. The dashed line indicates the performance of APS without SACP.
V-C Discussion
SACP is robust to the size of calibration set
As shown in Eq. (4), the calibration set is used to determine the threshold $\tau$ corresponding to the error rate of $1-\alpha$. A precise estimate of $\tau$ is essential. In this work, all experiments are conducted with a fixed ratio for the calibration set size, specifically $|\mathcal{D}_{\text{cal}}|\approx 0.5\times|\mathcal{D}_{\text{cal}}\cup\mathcal{D}_{\text{test}}|$. In this section, we investigate whether SACP is sensitive to the size of the calibration set $|\mathcal{D}_{\text{cal}}|$. Specifically, we exam whether Coverage and Size remain satisfactory when varying the ratio of the calibration set size. The experiment is conducted on all three datasets using SSTN, with APS as the base score function and the error rate $\alpha$ set to $0.05$. The ratio of calibration samples is defined as $\gamma=(|\mathcal{D}_{\text{cal}}|/|\mathcal{D}_{\text{cal}}\cup\mathcal{D}_{\text{test}}|)$. $\gamma$ ranges from 0.1 to 0.6 in increments of 0.05. The experimental results are presented in Table IV. Coverage consistently maintains $1-\alpha$ as the calibration set size decreases relative to $|\mathcal{D}_{\text{cal}}\cup\mathcal{D}_{\text{test}}|$, even for very small values of $\gamma$. Additionally, the size of the prediction sets remains stable, only marginally influenced by changes in the calibration set size, and fluctuates within a narrow range. These findings collectively demonstrate that SACP exhibits significant robustness to variations in the calibration set size.
TABLE IV: Effect of Calibration Set Size on SACP: The ratio of calibration samples is given by $\gamma=\frac{|\mathcal{D}_{\text{cal}}|}{|\mathcal{D}_{\text{cal}}\cup\mathcal{D}_{\text{test}}|}$.
$\mathcal{D}_{\mathcal{I}}$ $\gamma=0.1$ $\gamma=0.15$ $\gamma=0.2$ $\gamma=0.25$ $\gamma=0.3$ $\gamma=0.35$ $\gamma=0.4$ $\gamma=0.45$ $\gamma=0.5$ $\gamma=0.55$ $\gamma=0.6$
IP Coverage 0.96 0.95 0.95 0.95 0.95 0.95 0.95 0.95 0.95 0.95 0.95
Size 1.83 1.70 1.73 1.75 1.74 1.73 1.75 1.73 1.73 1.76 1.72
PU Coverage 0.95 0.95 0.95 0.95 0.95 0.95 0.95 0.95 0.95 0.95 0.95
Size 1.24 1.25 1.24 1.24 1.26 1.25 1.25 1.24 1.24 1.25 1.24
SA Coverage 0.96 0.95 0.95 0.95 0.95 0.95 0.95 0.95 0.95 0.95 0.95
Size 1.23 1.20 1.21 1.19 1.19 1.20 1.20 1.19 1.19 1.20 1.19
SACP Exhibits Adaptiveness
Adaptiveness is a crucial attribute of CP, ensuring that prediction sets accurately reflect the uncertainty associated with each instance. Specifically, prediction sets should be larger for more challenging instances and smaller for easier ones. This analysis evaluates whether SACP maintains this characteristic. We use the true label rank $o(y,\hat{\pi}_{\theta}(\bm{x}))$ to represent the difficulty of the instance $\bm{x}$, where a higher true label rank indicates greater classification difficulty. Instances are categorized into five groups based on their true label rank: $1$, $2$, $3$, $4$, and $\{5,\cdots,|\mathcal{Y}|\}$. We then compute the average size of prediction sets for each group. This experiment is conducted on the SA and PU datasets using SSTN, with APS as the base score function and the error rate $\alpha$ set to $0.05$. Figure 4 demonstrates that SACP exhibits adaptiveness across various datasets. Both APS, with or without SACP, consistently show that prediction sets are smaller for easier instances compared to more challenging ones. SACP effectively optimizes the average size of prediction sets for each difficulty group while maintaining its adaptive characteristics. These results collectively underscore the significance of SACP in HSI classification.
(a) 1
(b) 2
Figure 4: Size vs. Difficulty: The term “Difficulty” refers to the ranking of the true class label in the output probability, reflecting the difficulty of accurately classifying the pixel. We report the average size of prediction sets for the PU and SA datasets using SSTN, categorized by their Difficulty.
SACP is computationally efficient
In this section, we analyze the computational efficiency of SACP. For the base score functions APS and RAPS, the computation involves operations such as sorting and summation. Consequently, the time complexity for computing both basic scores is $\mathcal{O}(|\mathcal{Y}|\log|\mathcal{Y}|N)$, where $N$ denotes the number of instances and $|\mathcal{Y}|$ represents the total number of labels. The dominant computational overhead in this process arises from the sorting operation. SACP introduces an additional aggregation step, contributing an extra time complexity of $\mathcal{O}(kN)$, where $k$ is the number of iterations. Thus, the overall time complexity of SACP is $\mathcal{O}(|\mathcal{Y}|\log|\mathcal{Y}|N+kN)$. For simplicity, we approximate $(k+|\mathcal{Y}|\log|\mathcal{Y}|)$ as a constant $C$, since these terms are negligible compared to $N$. Thus, the complexity can be expressed as $\mathcal{O}(CN)$, which remains comparable in magnitude to that of standard CP. The above analysis demonstrates that SACP exhibits computational efficiency.
VI Conclusion
In this paper, we investigate the application of CP to HSI classification and introduce a conformal procedure that integrates conformal prediction sets into HSI classifiers, which ensure the inclusion of true labels with a user-specified probability. Building on this, we propose SACP, a framework designed to enhance the efficiency of CP for HSI classification. The key idea of SACP is its utilization of intrinsic spatial information in HSIs to refine non-conformity scores. To the best of our knowledge, this is the first study to explore CP within the context of HSI data. Extensive experiments reveal that SACP outperforms standard CP in HSI classification by reducing the average prediction set size while ensuring satisfactory marginal and conditional coverage rates. Moreover, SACP is practical and user-friendly, showing robustness to hyperparameters and incurring minimal computational cost.
Limitations. The number of pixels within different classes in HSIs can sometimes be imbalanced, leading to unusually large prediction sets and potential undercoverage for classes with fewer pixels. This issue has not been addressed in our work and represents a promising avenue for future research.
Appendix A Proof of Theorem 2
Proof.
Here, we emphasize that the non-conformity score $S(B,y)$ is dependent on the sample partitioning, i.e.,
$S(B,y)\equiv S(B,y;\mathcal{D}_{\text{train}},\mathcal{D}_{\text{val}},\mathcal{D}_{\text{cal}}\cup\mathcal{D}_{\text{test}}).$ (19)
By Assumption 1, the non-conformity score $S(B,y)$ remains invariant under any permutation $\pi$ of the union $\mathcal{D}_{\text{cal}}\cup\mathcal{D}_{\text{test}}$. Therefore, we have
$S(B,y;\mathcal{D}_{\text{train}},\mathcal{D}_{\text{val}},\mathcal{D}_{\text{cal}}\cup\mathcal{D}_{\text{test}})$ (20)
$=$ $S(B,y;\mathcal{D}_{\text{train}},\mathcal{D}_{\text{val}},(\mathcal{D}_{\text{cal}}\cup\mathcal{D}_{\text{test}})^{\pi}),$
where $(\mathcal{D}_{\text{cal}}\cup\mathcal{D}_{\text{test}})^{\pi}$ denotes permuted samples according to $\pi$, and $\pi$ is permutation of $|\mathcal{D}_{\text{cal}}\cup\mathcal{D}_{\text{test}}|$. Hence, the score $S(B,y)$ remains invariant with respect to the selection of any subset of $\mathcal{D}_{\text{cal}}\cup\mathcal{D}_{\text{test}}$ as the calibration set $\mathcal{D}_{\text{cal}}$. Let $[t_{1},t_{2},\ldots,t_{n+m}]$ be the unordered scores from $\{S(B_{i},y_{i})\}_{i\in\mathcal{D}_{\text{cal}}\cup\mathcal{D}_{\text{test}}}$, where []delimited-[][$\cdot$][ ⋅ ] denotes unordered sets. We have $\{s_{i}\}_{i=1}^{n}$ is a subset of size $n$ of $[t_{1},t_{2},\ldots,t_{n+m}]$. Under random splitting of $\mathcal{D}_{\text{cal}}\cup\mathcal{D}_{\text{test}}$, any permutation $\pi$ happens with the same probability, i.e.,
($\mathbb{P}\left(\{s_{i}\}_{i=1}^{n}=\{t_{1},t_{2},\ldots,t_{n}\}\mid\{s_{i}\}_{i=1}^{n+m}\right)=\frac{1}{\binom{n+m}{n}}.$ (21)
As a result, any score of test sample in $\{s_{n+j}\}_{j=1}^{m}$ is exchangeable with $\{s_{i}\}_{i=1}^{n}$. Then, we complete the proof. ∎
Appendix B Proof of Proposition 1
Proof.
We introduce the superscript $\alpha$ to $\tau$ as follows:
$\tau^{\alpha}=\inf\left\{s\mid\frac{\left|\{i\mid s_{i}\leq s\}\right|}{N_{2}}\geq\frac{\lceil(N_{2}+1)(1-\alpha)\rceil}{N_{2}}\right\}.$ (22)
Denote the cuboids in $\mathcal{D}_{\text{test}}$ as $\mathcal{B}_{\text{test}}$. Then, for each cuboid $B$ in $\mathcal{B}_{\text{test}}$, we have
$\int_{0}^{1}|\mathcal{C}_{1-\alpha}(B)|\mathrm{d}\alpha$ $=\int_{0}^{1}|\{y\in\mathcal{Y}\mid S(B,y)\leq\tau^{\alpha}\}|~{}~{}\mathrm{d}\alpha$ (23)
$=\int_{0}^{1}\sum_{y\in\mathcal{Y}}\mathds{1}_{\{S(B,y))\leq\tau^{\alpha}\}}~{}~{}\mathrm{d}\alpha.$
Moreover, the event $\{S(B,y)\leq\tau^{\alpha}\}$ in Eq. (23) is equivalent to the event that at least $\alpha$ percent of the sample in $\mathcal{D}_{\text{cal}}$ have scores larger than that of $(B,y)$. . Thus, we have
$\int_{0}^{1}\sum_{y\in\mathcal{Y}}\mathds{1}_{\{S(B,y))\leq\tau^{\alpha}\}}~{}~{}\mathrm{d}\alpha$ (24)
$=$ $\int_{0}^{1}\sum_{y\in\mathcal{Y}}\mathds{1}_{\{\sum_{(\tilde{B},\tilde{y})\in\mathcal{D}_{\text{cal}}}\mathds{1}_{\{S(\tilde{B},\tilde{y})\geq S(B,y)\}}\geq\lfloor(1+|\mathcal{D}_{\text{cal}}|)\alpha\rfloor\}}~{}~{}\mathrm{d}\alpha$
$=$ $\int_{0}^{1}\sum_{y\in\mathcal{Y}}\mathds{1}_{\{\frac{1+\sum_{(\tilde{B},\tilde{y})\in\mathcal{D}_{\text{cal}}}\mathds{1}_{\{S(\tilde{B},\tilde{y})\geq S(B,y)\}}}{1+|\mathcal{D}_{\text{cal}}|}>\alpha\}}~{}~{}\mathrm{d}\alpha.$
It is obvious that
$0<\frac{1+\sum_{(\tilde{B},\tilde{y})\in\mathcal{D}_{\text{cal}}}\mathds{1}_{\{S(\tilde{B},\tilde{y})\geq S(B,y)\}}}{1+|\mathcal{D}_{\text{cal}}|}\leq 1$ (25)
Thus, we have
$\int_{0}^{1}\sum_{y\in\mathcal{Y}}\mathds{1}_{\{\frac{1+\sum_{(\tilde{B},\tilde{y})\in\mathcal{D}_{\text{cal}}}\mathds{1}_{\{S(\tilde{B},\tilde{y})\geq S(B,y)\}}}{1+|\mathcal{D}_{\text{cal}}|}>\alpha\}}~{}~{}\mathrm{d}\alpha$ (26)
$=$ $\sum_{y\in\mathcal{Y}}\int_{0}^{1}\mathds{1}_{\{\frac{1+\sum_{(\tilde{B},\tilde{y})\in\mathcal{D}_{\text{cal}}}\mathds{1}_{\{S(\tilde{B},\tilde{y})\geq S(B,y)\}}}{1+|\mathcal{D}_{\text{cal}}|}>\alpha\}}~{}~{}\mathrm{d}\alpha$
$=$ $\sum_{y\in\mathcal{Y}}\frac{1+\sum_{(\tilde{B},\tilde{y})\in\mathcal{D}_{\text{cal}}}\mathds{1}_{\{S(\tilde{B},\tilde{y})\geq S(B,y)\}}}{1+|\mathcal{D}_{\text{cal}}|}$
For $\mathcal{B}_{\text{test}}$, we have
$\sum_{B\in\mathcal{B}_{\text{test}}}\int_{0}^{1}|\mathcal{C}_{1-\alpha}(B)|\,\mathrm{d}\alpha$ (27)
$=$ $\sum_{B\in\mathcal{B}_{\text{test}}}\sum_{y\in\mathcal{Y}}\frac{1+\sum_{(\tilde{B},\tilde{y})\in\mathcal{D}_{\text{cal}}}\mathds{1}_{\{S(\tilde{B},\tilde{y})\geq S(B,y)\}}}{1+|\mathcal{D}_{\text{cal}}|}$
$=$ $\sum_{(B,y)\in\hat{\mathcal{D}}_{\text{test}}}\frac{1+\sum_{(\tilde{B},\tilde{y})\in\mathcal{D}_{\text{cal}}}\mathds{1}_{\{S(\tilde{B},\tilde{y})\geq S(B,y)\}}}{1+|\mathcal{D}_{\text{cal}}|}$
$=$ $\sum_{(B,y)\in\hat{\mathcal{D}}_{\text{test}}}\frac{1+\sum_{(\tilde{B},\tilde{y})\in\mathcal{D}_{\text{cal}}}\mathds{1}_{\{S(\tilde{B},\tilde{y})\geq S(B,y)\}}}{1+n}$
$=$ $\frac{n^{\prime}|\mathcal{Y}|}{1+n}+\frac{1}{1+n}\sum_{i=1}^{n}\sum_{k=1}^{n^{\prime}|\mathcal{Y}|}\mathds{1}_{\{s_{i}>t_{k}\}}.$
In addition, we have
${R}(\{s_{i}\}_{i=1}^{n},\{t_{k}\}_{k=1}^{n^{\prime}|\mathcal{Y}|})=\frac{1}{n\times n^{\prime}|\mathcal{Y}|}\sum_{i=1}^{n}\sum_{k=1}^{n^{\prime}|\mathcal{Y}|}\mathds{1}_{\{s_{i}>t_{k}\}},$ (28)
Finally, we have
$\sum_{(B,y)\in\mathcal{D}_{\text{test}}}\int_{0}^{1}|\mathcal{C}_{1-\alpha}(B)|\,\mathrm{d}\alpha$ (29)
$=$ $\frac{n\times n^{\prime}|\mathcal{Y}|}{1+n}{R}(\{s_{i}\}_{i=1}^{n},\{t_{k}\}_{k=1}^{n^{\prime}|\mathcal{Y}|})+\frac{n^{\prime}|\mathcal{Y}|}{1+n},$
where $\frac{n\times n^{\prime}|\mathcal{Y}|}{1+n}$ and $\frac{n^{\prime}|\mathcal{Y}|}{1+n}$ are positive constants. Based on Eq. (29), we obviously have
${R}(\{s_{i}\}_{i=1}^{n},\{t_{k}\}_{k=1}^{n^{\prime}|\mathcal{Y}|})$ $>{R}(\{\hat{{s}}_{i}\}_{i=1}^{n},\{\hat{t}_{k}\}_{k=1}^{n^{\prime}|\mathcal{Y}|})\Leftrightarrow$ (30)
$\sum_{(B,y)\in\mathcal{D}_{\text{test}}}\int_{0}^{1}|\mathcal{C}_{1-\alpha}(B)|\,\mathrm{d}\alpha$ $>\sum_{(B,y)\in\mathcal{D}_{\text{test}}}\int_{0}^{1}|\hat{\mathcal{C}}_{1-\alpha}(B)|\,\mathrm{d}\alpha.$
Then, we complete the proof. ∎
Appendix C Proof of Proposition 2
Proof.
We will employ mathematical induction to demonstrate that the statement is valid for all $k$.
Base Case ($k=0$)
For $k=0$, the operator $\mathcal{V}_{k}$ reduces to $S$. By our Theorem 2, we have $\hat{s}_{i},(i=1,\cdots,n)$ and $S(B_{n+1},y_{n+1})$ are exchangeable. Thus,
$\mathbb{P}\left(y_{n+1}\notin\hat{\mathcal{C}}_{1-\alpha}(B_{n+1};\hat{\tau})\right)$ (31)
$=$ $\mathbb{P}\left(S(B_{n+1},y_{n+1})>\hat{\tau}\right)$
$=$ $\frac{1}{n+1}\sum_{i=1}^{n+1}\mathbb{P}\left(S(B_{i},y_{i})>\hat{\tau}\right)$
$=$ $\frac{1}{n+1}\sum_{i=1}^{n+1}\mathbb{E}\mathds{1}_{\{i:S(B_{i},y_{i})>\hat{\tau}\}}$
$=$ $\frac{1}{n+1}\mathbb{E}|\{i:S(B_{i},y_{i})>\hat{\tau}\}|$
$\leq$ $\frac{1}{n+1}(n+1)\alpha$
$=$ $\alpha.$
Therefore, we have
$\mathbb{P}\left(y_{n+1}\in\hat{\mathcal{C}}_{1-\alpha}(B_{n+1};\hat{\tau})\right)\geq 1-\alpha$ (32)
Inductive Hypothesis
Assume the proposition holds for $k=m$. Specifically, assume:
$\mathbb{P}\left(y_{n+1}\in\hat{\mathcal{C}}_{1-\alpha}(B_{n+1};\hat{\tau})\right)\geq 1-\alpha,$ (33)
where
$\hat{\mathcal{C}}_{1-\alpha}(B_{n+1};\hat{\tau}):=\{y\in\mathcal{Y}\mid\mathcal{V}_{m}(B_{n+1},y)\leq\hat{\tau}\},$ (34)
and $\hat{s}_{i}^{(m)},i=1,\dots,n+1$ are exchangeable.
Inductive Step
To prove the proposition for $k=m+1$, we need to show:
$\mathbb{P}\left(y_{n+1}\in\hat{\mathcal{C}}_{1-\alpha}(B_{n+1};\hat{\tau})\right)\geq 1-\alpha,$ (35)
where
$\hat{\mathcal{C}}_{1-\alpha}(B_{n+1};\hat{\tau}):=\{y\in\mathcal{Y}\mid\mathcal{V}_{m+1}(B_{n+1},y)\leq\hat{\tau}\},$ (36)
and $\hat{s}_{i}^{(m+1)},i=1,\dots,n+1$ are exchangeable.
The aggregation operator for $k=m+1$ is:
$\mathcal{V}_{m+1}(B_{i},y)=(1-\lambda)\mathcal{V}_{m}(B_{i},y)+\frac{\lambda}{|\mathcal{N}_{i}|}\sum_{B_{j}\in\mathcal{N}_{i}}\mathcal{V}_{m}(B_{j},y),$ (37)
[…]
Then, we complete the proof. ∎